# AI-Generated Compromises for Coalition Formation: Modeling, Simulation, and a Textual Case Study

### v1

@@ 5.1 Interpretation
Our simulations demonstrate the effectiveness of AI-mediated coalition formation, particularly when leveraging Large Language Models (LLMs). AI-mediation significantly reduces the number of iterations required for coalitions to reach a compromise while minimizing the average distance between the final compromise and individual agents’ ideal documents. Notably, LLMs combined with distance-based optimization consistently accelerate convergence compared to simpler AI-mediator approaches.
The statistical tests reinforce that meaningful differences exist among mediation strategies, underscoring the adaptability of AI-mediation to different scenarios. The superior performance of Option 2 in minimizing iterations further suggests that careful tuning of the AI-mediator’s behavior can yield substantial efficiency gains.
@@ 5.2 Future Work
We outline several directions to extend the current model:
@@
  • Core-Like Stability: Introduce blocking coalitions: subsets of agents that can jointly deviate to strictly preferred compromise points. This enables core-inspired stability notions.
@@ A.1 A Concrete Example
    Iteration 1: The AI-mediator suggests the compromise point $p^{BC}$ to the coalitions $(C_{B},p^{B})$ and $(C_{C},p^{C})$. Both agents approve since $1<6$ and $1<8$. We arrive to the following coalition structure $D^{\prime}$:
@@
A.2 AI-Mediators in a 2D Euclidean Space
@@ A.2.1 Simulation-Based Analysis
  • Status Quo: Generated uniformly at random between $(0,200)\times(0,200)$.
  • Ideal Points: Drawing inspiration from the literature [9], each agent was assigned an ideal point $(x,y)$ with both coordinates sampled from either the uniform distribution between $(0,200)$ or from a 2-dimensional Gaussian Mixture Model (GMM). The GMM represents the overall probability distribution as a weighted sum of several Gaussian components with multiple peaks. In our simulations, we considered GMMs with $g$ combined Gaussian distributions, for $g\in\{0,1,2,3,4\}$ with the mean of each Gaussian being distributed uniformly between 00 and $200$ in each dimension, its deviation distributed uniformly between 00 and $50$, and the weights signifying the importance of each Gaussian are distributed from the $Dirichlet(\alpha^{g}\in{\mathbb{R}_{0}^{+}}^{g})$ distribution with $\alpha$ set to $1$ (resulting in $g$ numbers that sum to $1$). This sampling of ideal points is demonstrated in the supplementary material. Note that we treat GMM with $g=0$ (i.e., 00 peaks) as the uniform distribution.
@@
  • For the AI-mediator, we have used $\alpha=\{-1,0,1\}$ (as described in realization Subsection int he main text).
@@
We conducted $100$ independent repetitions for each configuration. Next we present our two evaluation metrics (the first measures the process speed, while the second measures the process quality):
@@
For efficiency, we halt our simulations whenever the number of iterations exceeds a threshold of $10,000$ (i.e., we treat an instance for which no convergence is reached within $10,000$ iterations as an instance that does not converge at all).
@@ A.2.2 Results and Discussion
Next we discuss the main conclusions, drawn at a $5\%$ significance level:

### v2

@@ 5.1 Interpretation
Our simulations demonstrate the effectiveness of AI-mediated coalition formation, particularly when leveraging Large Language Models (LLMs). AI-mediation significantly reduces the number of iterations required for coalitions to reach a compromise while minimizing the average distance between the final compromise and individual agents’ ideal documents. Notably, LLMs combined with distance-based optimization consistently accelerate convergence compared to simpler mediator approaches.
The statistical tests reinforce that meaningful differences exist among mediation strategies, underscoring the adaptability of AI-mediation to different scenarios. The superior performance of Option 2 in minimizing iterations further suggests that careful tuning of the mediator’s behavior can yield substantial efficiency gains.
@@ 5.2 Future Work
Several directions for future research are outlined below:
@@
  • Strategic Behavior: Extend the model to a game-theoretic setting that accounts for strategic agents who may misreport preferences and anticipate coalition dynamics.
@@
  • Context-Sensitive Voting Behavior: Our model compares proposals to a fixed status quo, following Elkind et al. [10]. A natural extension is to compare proposals to the current coalition point $p_{c}$, i.e., accept $p$ if $d(p^{v},p)<d(p^{v},p_{c})$. A more refined model would also consider coalition size and composition—agents may prefer large coalitions for influence, or avoid those with ideologically distant members.
@@ A.1 A Concrete Example
    Iteration 1: The mediator suggests the compromise point $p^{BC}$ to the coalitions $(C_{B},p^{B})$ and $(C_{C},p^{C})$. Both agents approve since 1<61<61 < 6 and 1<81<81 < 8. We arrive to the following coalition structure $D^{\prime}$:
@@
A.2 Mediators in a 2D Euclidean Space
@@ A.2.1 Simulation-Based Analysis
  • Status Quo: Generated uniformly at random between $\times(0,200)$ ( 0 , 200 ).
  • Ideal Points: Drawing inspiration from the literature [9], each agent was assigned an ideal point $(x,y)$ with both coordinates sampled from either the uniform distribution between (0,200)(0,200)( 0 , 200 ) or from a 2-dimensional Gaussian Mixture Model (GMM). The GMM represents the overall probability distribution as a weighted sum of several Gaussian components with multiple peaks. In our simulations, we considered GMMs with $g$ combined Gaussian distributions, for $g\in\{0,1,2,3,4\}$ with the mean of each Gaussian being distributed uniformly between 0 and 200200200 in each dimension, its deviation distributed uniformly between 0 and 505050, and the weights signifying the importance of each Gaussian are distributed from the $Dirichlet(\alpha^{g}\in{\mathbb{R}_{0}^{+}}^{g})$ distribution with $\alpha$ set to 111 (resulting in $g$ numbers that sum to 111). This sampling of ideal points is demonstrated in the supplementary material. Note that we treat GMM with $g=0$ (i.e., 0 peaks) as the uniform distribution.
@@
  • For the mediator, we have used $\alpha=\{-1,0,1\}$ (as described in realization Subsection int he main text).
@@
We conducted 100100100 independent repetitions for each configuration. Next we present our two evaluation metrics (the first measures the process speed, while the second measures the process quality):
@@
For efficiency, we halt our simulations whenever the number of iterations exceeds a threshold of 10,00010,00010 , 000 (i.e., we treat an instance for which no convergence is reached within 10,00010,00010 , 000 iterations as an instance that does not converge at all).
@@ A.2.2 Results and Discussion
[…]